OpenStudy (anonymous):

find the eigenvalues for the system of equations. x'(t) = [-7 0 0 0 8 -3 4 0 1 0 -5 0 2 1 4 -1 ] X(t) Thanks in advance!

OpenStudy (anonymous):



OpenStudy (theeric):

Would this be like \(A=\begin{bmatrix}-7&0&0&0\\8&-3&4&0\\1&0&-5&0\\2&1&4&-1\end{bmatrix}\) and \(A\vec x=\lambda\vec x\) where all \(\lambda\) that make this true are called eigenvalues? I guess that \(x'(t)=\lambda\vec x(t)\), but I haven't seen the notation before. What you want to do is find the determinant \(\det(\lambda I-A)=0\) That's because \(A\vec x=\lambda\vec x\) \(\implies \lambda\vec x-A\vec x=0\) \(\implies (\lambda I-A)\vec x=0\) which has a non-zero solution for \(\vec x\) if the determinant of the multiplying matrix is not zero. If you need more help, I think there is a linear algebra group on OpenStudy! This is the ODE section. Take care!

OpenStudy (theeric):

When you find the determinant, it will have \(\lambda\) in it. But you set the determinant to zero and solve for \(\lambda\).